Can someone help me here?
I need to solve the system by addition or substitution
3x - 4y = 8
6x - 2y = 10
Thanks,
Jamie

Answer by stanbon(75887)

1st: 3x - 4y = 8
2nd: 6x - 2y = 10
------------
Multiply 1st by 2 to get:
3rd: 6x-8y=16
--------------
Subtract 2nd from 3rd to get:
-6y=6
y=-1
---------
Substitute into 1st so solve for "x":
3x-4(-1)=8
3x=4
x=4/3
----------
Solution:
x=4/3; y=-1
Cheers,
Stan H.